# Improving Image Classification Accuracy through Complementary Intra-Class and Inter-Class Mixup

### 21 Mar 2024 v1

@@ 4.1.1 Training Setup
In our experiments, we have meticulously curated a diverse experimental setup to effectively assess the robustness and generalizability of our integrated solution. This diversity is reflected in the selection of datasets, model input sizes, model architectures, and the use of pre-trained models.
@@
  • Model Input Size Diversity: Our selection of datasets incorporates a variety of model input sizes, from CIFAR-100’s $32\times 32$ pixels to the $224\times 224$ pixels of Food-101, Caltech-256, and OxfordIIIPet, with mini-ImageNet at an intermediate $84\times 84$ pixels. This variation tests the performance of our integrated solution across different input sizes.
  • Model Architecture Diversity: We have selected a range of model architectures, including the less complex ResNet-18 [32], the moderately complex ResNet-34 and ResNet-50 [32], the deeper ResNet-101 [32], the efficient MobileNetV3-Large [33], and the compact but advanced tiny-SwinTransformer [34]. This lineup allows us to evaluate our integrated solution’s adaptability across models with varying complexities and depths.
  • Utilization of Pre-trained Models: Our study investigates whether our method is effective when models are trained from scratch as well as when they are initialized with weights from pre-trained models. This is to validate that our integrated solution is robust and functions well under both conditions.
@@ 4.1.2 Comparable Methods
  • w-ER applies ER. ER is implemented by MixUp or Manifold MixUp, denoted by the suffix “(M)” or “(MM)”, respectively. The associated loss function is detailed in Eq. (3) or Eq. (4).
  • w-RA incorporates RA to the feature representations of unaugmented original images, with the loss function given by Eq. (8).
  • w-RA(aug) applies RA to the feature representations of augmented images. It’s loss function is detailed in Eq. (7).
  • w-RA&ER combines RA and ER, representing the culmination of our integrated solution. The corresponding loss function is Eq. (5).
@@ 4.1.3 Valiation Setup and Evaluation Metrics
During the validation stage, we allocated 10% of the training data from each dataset to form a validation set through stratified sampling, ensuring that the class distribution in the validation set mirrors that of the original training data. The remaining 90% of the data was used for training purposes. The validation set played a crucial role in fine-tuning the hyperparameter $\beta$. For w-RA, $\beta$ options were [0.2, 0.4, 0.5, 0.7], and for w-RA&ER(M) and w-RA&ER(MM), choices were [0.05, 0.1, 0.2, 0.4]. In the testing stage, the entire training set is used to train models using the best $\beta$ from the validation stage. Given that each dataset’s test set is balanced, we used accuracy as the sole metric for performance evaluation. To mitigate the effects of variability in model accuracy over epochs, we averaged the last 10 epochs for CIFAR-100 and Food-101 over 120 epochs, and the last 5 epochs for other datasets over 60 epochs. The effectiveness of our integrated solution was evaluated in achieving higher classification accuracy compared to when w-RA and w-ER are used independently.
@@
4.2 Performance Evaluation: Training from Scratch
In this subsection, we present and analysis the experimental results of the ResNet models of different depths trained from scratch. Here, our analytical approach begins with an exploration of the performance of w-RA, followed by an analysis of w-ER(M) and w-ER(MM). Subsequently, we delve into the performance of the intergated methods, w-RA&ER(M) and w-RA&ER(MM), to ascertain their combined effects.
@@ 4.2 Performance Evaluation: Training from Scratch
4.2.1 Performance of w-RA
As shown in Figure. 5, an improvement in accuracy was observed across various mini-batch sizes, datasets, and depths of ResNet models when compared to the baseline method, wo-RA&ER. Specifically, w-RA outperformed wo-RA&ER in 9 out of 12 experimental scenarios, with accuracy gains ranging from about 0.22% to 3.09%, and an average increase of about 1.15%. In Figure , w-ER(M), w-ER(MM), and wo-RA&ER(model=ResNet-34) exhibited anomalously low classification accuracies due to a lack of stable convergence, whereas w-RA remained stable in convergence and achieved high classification accuracies with low standard deviations. These results exhibit that applying the mixing operation to feature representations of unaugmented original images within the same class has the notable potential to improve classification accuracy and robustness to model convergence.
@@ 4.2.1 Performance of w-RA
It is worthing note that there are three experimental scenarios where w-RA underperforms wo-RA&ER in Figure . This could be attributed to several factors. Firstly, the validation set, comprising a random 10% subset of the train set, may have a distribution that differs from that of the test set, potentially impacting the selection of the optimal $\beta$ value. This effect was particularly pronounced in the Food-101 dataset, where the training set contains noisy data and incorrect labels, in contrast to the clean test set [6]. Secondly, since there is no guarantee that the synthesized feature representations generated by w-RA are sampled from the true distribution, they can mislead the classifier, making it more difficult to correctly distinguish between different classes. This is particularly problematic if the synthesized feature representations of one class overlap with the feature distribution of another class. It increases the likelihood of misclassification and thus reduce the overall accuracy of the model. Notably, this problem can be alleviated by combining with w-ER, as illustrated in Section 4.2.3.
@@ 4.2.2 Performance of w-ER
Figure  shows that w-ER(M) achieved accuracy gains compared to wo-RA&ER in 11 out of 12 experimental scenarios, with improvements ranging from about 0.54% to 4.35%, and an average increase of 1.92%. w-ER(MM) also saw accuracy gains in 11 out of 12 experimental scenarios, with increases ranging from about 1% to 3.47%, and an average of 2.32%. Furthermore, it also achieved higher accuracy than w-ER(M) in 10 out of the 12 scenarios, by approximately 0.31% to 1.64%, averaging 0.97%. The performance of w-ER(M) and w-ER(MM) is consistent with the results reported in the literature [1, 2].
@@
4.2.3 Performance of w-RA&ER
In every experimental setup of Figure , w-RA&ER outperformed the individual use of w-RA and w-ER. Specifically, w-RA&ER(M) achieved an average accuracy increase of 2.61%, with gains ranging from 0.78% to 3.95% over w-RA. Against w-ER(M), it improved by an average of 1.66%, with a range of 0.58% to 4.29%. Especially, a synergistic effect was observed in 8 of 12 experimental scenarios such as with the CIFAR-100 dataset using a ResNet-18 model at a mini-batch size of 32, where w-RA&ER(M) exceeded the sum of the individual gains achieved by w-RA and w-ER(M) alone. For w-RA&ER(MM), the improvements were also significant. It surpassed w-RA with an average increase of 3.05%, and gains varied from 1.46% to 5.16%. When compared to w-ER(MM), the average increase was 1.97%, with a range of 0.12% to 11.77%. A synergistic effect was also noted in 8 out of the 12 scenarios. Collectively, these findings demonstrate that the benefits (as stated in Section 3.2’s third question) of intra-class mixup and inter-class mixup are successfully integrated, significantly enhancing classification accuracy.
@@ 4.2.3 Performance of w-RA&ER
Although w-RA and w-ER complement each other, their combination w-RA&ER does not guarantee improved accuracy in all scenarios. As shown in Figure , for the ResNet-34 model, w-RA&ER(M) and w-RA&ER(MM) demonstrated notable accuracy gains over w-ER(M) and w-ER(MM), but they did not surpass the accuracy of w-RA alone. This pattern was also observed with w-RA&ER(M) in the ResNet-18 model. To address this issue, increasing the mini-batch size can enhance the convergence stability of w-ER, potentially maximizing the benefits derived from their combination.
@@ 4.3 Performance Evaluation: Fine-Tuning using Pre-Trained Model
In this subsection, we report the experimental results of deep learning models with different architectures that were pretrained on the ImageNet-1K dataset. As shown in Figure 9, w-RA&ER still achieved higher classification accuracy than when w-RA and w-ER were used individually. When these findings are integrated with those from Section 4.2, it becomes evident that our integrated solution is versatile. It proves effective not only for training strategies that begin from scratch but also for fine-tuning pre-trained models. Furthermore, it is adaptable across models of different depths and architectures and is capable of handling image classification tasks with a wide range of classes.
@@ 4.4.1 Investigation of the Hyperparameter $\beta$
$\beta$ is the only new hyperparameter introduced in our integrated solution, serving to balance the benefits of improving intra-class cohesion against the benefits of augmenting inter-class distinction. The choice of this parameter’s value is critical for the model’s performance. To explore the impact of this parameter on model performance, we conducted extensive experiments depicted in Figure 10, which considered variations in model depth, mini-batch size, and dataset.
@@ 4.4.2 Investigation of the Number $P$ of Linear Interpolations
As stated in Section 3.2’s second question, our intra-class mixup method prefers the use of feature representations from unaugmented original images over those from augmented images to create synthesized feature representations. The soundness of this practice is corroborated by Figure 12. Figures 12(a) and 12(b) both illustrate that synthesized feature representations derived from feature representations of unaugmented original images yield superior classification accuracy across two distinct $\beta$ values compared to those derived from augmented images.
@@ 4.5 Discussion
Our integrated solution incorporates both inter-class mixup happened at the input or early hidden layers and intra-class mixup occurred at the penultimate layer before the single fully connected layer dedicated to classification. These layers are present in a variety of model architectures, making our solution largely agnostic to the specific model architecture. As a result, it can be easily transferred to other domains where current mixup techniques have demonstrated potential, such as in speech and text classification. Additionally, our solution can be tailored for tasks involving classification components, such as object detection.
Despite the notable gains of our solution in classification accuracy, it also leads to an almost twofold increase in training costs. This is attributed to our intra-class mixup method, which adds a branch that closely mirrors the original model. For example, on the Food-101 dataset, using a ResNet-101 model with a mini-batch size of 32 on an NVIDIA GTX 4090 graphics card, the epoch duration of wo-RA&ER is roughly 290 seconds. In contrast, the training time of w-RA&ER(M) extends to about 570 seconds per epoch.
Potential avenues for future research include: (1) Our intra-class mixup method has proven effective in markedly accelerating model convergence, as measured by epoch count. Building upon this, further research could analyze the shifting trajectories of feature representations across epochs. Using these trajectories, predictive synthesized feature representations could be generated to further expedite model convergence, mitigating the increased computational cost problem mentioned above. (2) As illustrated in Section 4.4.2, creating synthesized feature representations within the convex hull or on the line segment introduces beneficial stochasticity into the gradients, potentially contributing to higher accuracy. This observation encourages more in-depth research into the impact of the randomness of synthesized representations within local feature spaces on gradient stochasticity in order to improve classification accuracy.
@@ 5 Conclusion
Existing mixup techniques often overlook intra-class mixup in terms of mixup strategies (intra-class and inter-class mixup), and fall short in enhancing intra-class cohesion through their mixing operations in terms of mixing outcomes. These limitations constrain their performance in image classification tasks. In response to these shortcomings, this article has undertaken the following work:
(1) We have introduced a novel intra-class mixup method that leverages feature representations of unaugmented original images from the same class to generate synthesized feature representations, with the aim of bolstering intra-class cohesion. The principle behind the method, along with specific implementation strategies—such as giving preference to feature representations from unaugmented original images over those from augmented images for synthesis, and restricting the generation to one synthesized feature representation per class per batch—has been thoroughly investigated. Our empirical research supports that, compared to the baseline devoid of any mixup operations, our intra-class mixup method reduces intra-class variability and boosts classification accuracy by an average of 1.25% (ranging from 0.22% to 3.09%) in 15 out of 18 experimental scenarios. Moreover, it shows considerable promise in markedly accelerating model convergence, as indicated by a reduction in the total number of training epochs required.
(2) Building upon our intra-class mixup method and established inter-class mixup methods such as MixUp and Manifold Mixup, we have proposed a comprehensive integrated solution This solution incorporates inter- and intra-class mixup in a balanced manner while concurrently enhancing intra-class cohesion and inter-class separability, effectively overcoming the two limitations identified in existing mixup methods. Empirical results from a variety of experimental scenarios indicate that our solution outperforms the best of either MixUp or our intra-class mixup method in accuracy by margins ranging from 0.1% to 3.43% (averaging 1.16%). It also surpasses the superior of either Manifold MixUp or our intra-class mixup method by margins ranging from 0.12% to 5.16% (averaging 1.11%).
Our integrated solution is versatile and can be applied to other research areas and tasks, such as speech classification, text classification, and object detection. It also paves the way for new potential research directions, which include: a) developing methods for predictive synthesis of feature representations, which leverage the shifting trajectories of image feature representations across epochs to accelerate model convergence, and b) investigating how the randomness of synthesized feature representations within local feature spaces affects the stochasticity of the gradient, with the aim of enhancing classification accuracy.

### 24 Mar 2024 v2

@@ 4.1.1 Training Setup
In our experiments, we have meticulously curated a diverse experimental setup to effectively assess the robustness and generalizability of SynerMix. This diversity is reflected in the selection of datasets, model input sizes, model architectures, and the use of pre-trained models.
@@
  • Model Input Size Diversity: Our selection of datasets incorporates a variety of model input sizes, from CIFAR-100’s $32\times 32$ pixels to the $224\times 224$ pixels of Food-101, Caltech-256, and OxfordIIIPet, with mini-ImageNet at an intermediate $84\times 84$ pixels. This variation tests the performance of SynerMix across different input sizes.
  • Model Architecture Diversity: We have selected a range of model architectures, including the less complex ResNet-18 [32], the moderately complex ResNet-34 and ResNet-50 [32], the deeper ResNet-101 [32], the efficient MobileNetV3-Large [33], and the compact but advanced tiny-SwinTransformer [34]. This lineup allows us to evaluate SynerMix’s adaptability across models with varying complexities and depths.
  • Utilization of Pre-trained Models: Our study investigates whether our method is effective when models are trained from scratch as well as when they are initialized with weights from pre-trained models. This is to validate that SynerMix is robust and functions well under both conditions.
@@ 4.1.2 Comparable Methods
  • w-ER only applies inter-class mixup. This method is implemented through MixUp or Manifold MixUp, denoted by the suffix “(M)” or “(MM)”, respectively. Its associated loss functions are detailed in Eq. (3) or Eq. (4) accordingly.
  • w-RA represents our proposed mixup method SynerMix-Intra. It performs intra-class mixup on the feature representations of unaugmented original images, with the loss function given by Eq. (8).
  • w-RA(aug) applies intra-class mixup to the feature representations of augmented images. The corresponding loss function is detailed in Eq. (7).
  • w-RA&ER indicates our synergistic mixup solution SynerMix, which combines w-RA and w-ER. Its loss function is Eq. (5).
@@ 4.1.3 Valiation Setup and Evaluation Metrics
During the validation stage, we allocated 10% of the training data from each dataset to form a validation set through stratified sampling, ensuring that the class distribution in the validation set mirrors that of the original training data. The remaining 90% of the data was used for training purposes. The validation set played a crucial role in fine-tuning the hyperparameter $\beta$. For w-RA, $\beta$ options were [0.2, 0.4, 0.5, 0.7], and for w-RA&ER(M) and w-RA&ER(MM), choices were [0.05, 0.1, 0.2, 0.4]. In the testing stage, the entire training set is used to train models using the best $\beta$ from the validation stage. Given that each dataset’s test set is balanced, we used accuracy as the sole metric for performance evaluation. To mitigate the effects of variability in model accuracy over epochs, we averaged the last 10 epochs for CIFAR-100 and Food-101 over 120 epochs, and the last 5 epochs for other datasets over 60 epochs. The effectiveness of SynerMix was evaluated in achieving higher classification accuracy compared to when w-RA and w-ER are used independently.
@@
4.2 Performance Evaluation: Training from Scratch
In this subsection, we present and analysis the experimental results of the ResNet models of different depths trained from scratch. Here, our analytical approach begins with an exploration of the performance of w-RA, followed by an analysis of w-ER(M) and w-ER(MM). Subsequently, we delve into the performance of the synergistic mixup solutions, w-RA&ER(M) and w-RA&ER(MM), to ascertain their combined effects.
@@ 4.2 Performance Evaluation: Training from Scratch
4.2.1 Performance of w-RA (SynerMix-Intra)
As shown in Figure 5, an improvement in accuracy was observed across various mini-batch sizes, datasets, and depths of ResNet models when compared to the baseline method, wo-RA&ER. Specifically, w-RA, namely SynerMix-Intra, outperformed wo-RA&ER in 9 out of 12 experimental scenarios, with accuracy gains ranging from about 0.22% to 3.09%, and an average increase of about 1.15%. In Figure 6, w-ER(M), w-ER(MM), and wo-RA&ER(model=ResNet-34) exhibited anomalously low classification accuracies due to a lack of stable convergence, whereas w-RA remained stable in convergence and achieved high classification accuracies with low standard deviations. These results exhibit that applying the mixing operation to feature representations of unaugmented original images within the same class has the notable potential to improve classification accuracy and robustness to model convergence.
@@ 4.2.1 Performance of w-RA (SynerMix-Intra)
It is worthing note that there are three experimental scenarios where w-RA underperforms wo-RA&ER in Figure 5. This could be attributed to several factors. Firstly, the validation set, comprising a random 10% subset of the train set, may have a distribution that differs from that of the test set, potentially impacting the selection of the optimal $\beta$ value. This effect was particularly pronounced in the Food-101 dataset, where the training set contains noisy data and incorrect labels, in contrast to the clean test set [6]. Secondly, since there is no guarantee that the synthesized feature representations generated by w-RA are sampled from the true distribution, they can mislead the classifier, making it more difficult to correctly distinguish between different classes. This is particularly problematic if the synthesized feature representations of one class overlap with the feature distribution of another class. It increases the likelihood of misclassification and thus reduce the overall accuracy of the model. Notably, this problem can be alleviated by combining with w-ER, as illustrated in Section 4.2.3.
@@ 4.2.2 Performance of w-ER
Figure 5 shows that w-ER(M) achieved accuracy gains compared to wo-RA&ER in 11 out of 12 experimental scenarios, with improvements ranging from about 0.54% to 4.35%, and an average increase of 1.92%. w-ER(MM) also saw accuracy gains in 11 out of 12 experimental scenarios, with increases ranging from about 1% to 3.47%, and an average of 2.32%. Furthermore, it also achieved higher accuracy than w-ER(M) in 10 out of the 12 scenarios, by approximately 0.31% to 1.64%, averaging 0.97%. The performance of w-ER(M) and w-ER(MM) is consistent with the results reported in the literature [1, 2].
@@
4.2.3 Performance of w-RA&ER (SynerMix)
In every experimental setup of Figure 5, w-RA&ER, namely SynerMix, outperformed the individual use of w-RA and w-ER. Specifically, w-RA&ER(M) achieved an average accuracy increase of 2.61%, with gains ranging from 0.78% to 3.95% over w-RA. Against w-ER(M), it improved by an average of 1.66%, with a range of 0.58% to 4.29%. Especially, a synergistic effect was observed in 8 of 12 experimental scenarios such as with the CIFAR-100 dataset using a ResNet-18 model at a mini-batch size of 32, where w-RA&ER(M) exceeded the sum of the individual gains achieved by w-RA and w-ER(M) alone. For w-RA&ER(MM), the improvements were also significant. It surpassed w-RA with an average increase of 3.05%, and gains varied from 1.46% to 5.16%. When compared to w-ER(MM), the average increase was 1.97%, with a range of 0.12% to 11.77%. A synergistic effect was also noted in 8 out of the 12 scenarios. Collectively, these findings demonstrate that the benefits (as stated in Section 3.2’s third question) of intra-class mixup and inter-class mixup are successfully integrated, significantly enhancing classification accuracy.
@@ 4.2.3 Performance of w-RA&ER (SynerMix)
Although w-RA and w-ER complement each other, their combination w-RA&ER does not guarantee improved accuracy in all scenarios. As shown in Figure 6, for the ResNet-34 model, w-RA&ER(M) and w-RA&ER(MM) demonstrated notable accuracy gains over w-ER(M) and w-ER(MM), but they did not surpass the accuracy of w-RA alone. This pattern was also observed with w-RA&ER(M) in the ResNet-18 model. To address this issue, increasing the mini-batch size can enhance the convergence stability of w-ER, potentially maximizing the benefits derived from their combination.
@@ 4.3 Performance Evaluation: Fine-Tuning using Pre-Trained Model
In this subsection, we report the experimental results of deep learning models with different architectures that were pretrained on the ImageNet-1K dataset. As shown in Figure 9, w-RA&ER still achieved higher classification accuracy than when w-RA and w-ER were used individually. When these findings are integrated with those from Section 4.2, it becomes evident that w-RA&ER, namely SynerMix, is versatile. It proves effective not only for training strategies that begin from scratch but also for fine-tuning pre-trained models. Furthermore, it is adaptable across models of different depths and architectures and is capable of handling image classification tasks with a wide range of classes.
@@ 4.4.1 Investigation of the Hyperparameter $\beta$
$\beta$ is the only new hyperparameter introduced in SynerMix, serving to balance the benefits of improving intra-class cohesion against the benefits of augmenting inter-class distinction. The choice of this parameter’s value is critical for the model’s performance. To explore the impact of this parameter on model performance, we conducted extensive experiments depicted in Figure 10, which considered variations in model depth, mini-batch size, and dataset.
@@ 4.4.2 Investigation of the Number $P$ of Linear Interpolations
As stated in Section 3.2’s second question, SynerMix-Intra prefers the use of feature representations from unaugmented original images over those from augmented images to create synthesized feature representations. The soundness of this practice is corroborated by Figure 12. Figures 12(a) and 12(b) both illustrate that synthesized feature representations derived from feature representations of unaugmented original images yield superior classification accuracy across two distinct $\beta$ values compared to those derived from augmented images.
@@ 4.5 Discussion
SynerMix incorporates both inter-class mixup happened at the input or early hidden layers and intra-class mixup occurred at the penultimate layer before the single fully connected layer dedicated to classification. These layers are present in a variety of model architectures, making SynerMix largely agnostic to the specific model architecture. As a result, it can be easily transferred to other domains where current mixup techniques have demonstrated potential, such as in speech and text classification. Additionally, SynerMix can be tailored for tasks involving classification components, such as object detection.
Despite the notable gains of SynerMix in classification accuracy, it also leads to an almost twofold increase in training costs. This is attributed to SynerMix-Intra, which adds a branch that closely mirrors the original model. For example, on the Food-101 dataset, using a ResNet-101 model with a mini-batch size of 32 on an NVIDIA GTX 4090 graphics card, the epoch duration of wo-RA&ER is roughly 290 seconds. In contrast, the training time of w-RA&ER(M) extends to about 570 seconds per epoch.
Potential avenues for future research include: (1) SynerMix-Intra has proven effective in markedly accelerating model convergence, as measured by epoch count. Building upon this, further research could analyze the shifting trajectories of feature representations across epochs. Using these trajectories, predictive synthesized feature representations could be generated to further expedite model convergence, mitigating the increased computational cost problem mentioned above. (2) As illustrated in Section 4.4.2, creating synthesized feature representations within the convex hull or on the line segment introduces beneficial stochasticity into the gradients, potentially contributing to higher accuracy. This observation encourages more in-depth research into the impact of the randomness of synthesized representations within local feature spaces on gradient stochasticity in order to improve classification accuracy.
@@ 5 Conclusion
Existing mixup approaches often overlook intra-class mixup in terms of mixup strategies (intra-class and inter-class mixup), and fall short in enhancing intra-class cohesion through their mixing operations in terms of mixing outcomes. These limitations constrain their performance in image classification tasks. In response to these shortcomings, this article has undertaken the following work:
(1) We have introduced the novel mixup method SynerMix-Intra that leverages feature representations of unaugmented original images from the same class to generate synthesized feature representations, with the aim of bolstering intra-class cohesion. The principle behind the method, along with specific implementation strategies—such as giving preference to feature representations from unaugmented original images over those from augmented images for synthesis, and restricting the generation to one synthesized feature representation per class per batch—has been thoroughly investigated. Our empirical research supports that, compared to the baseline devoid of any mixup operations, SynerMix-Intra reduces intra-class variability and boosts classification accuracy by an average of 1.25% (ranging from 0.22% to 3.09%) in 15 out of 18 experimental scenarios. Moreover, it shows considerable promise in markedly accelerating model convergence, as indicated by a reduction in the total number of training epochs required.
(2) Building upon SynerMix-Intra and established inter-class mixup methods such as MixUp and Manifold Mixup, we have proposed the synergistic mixup solution SynerMix. SynerMix incorporates inter- and intra-class mixup in a balanced manner while concurrently enhancing intra-class cohesion and inter-class separability, effectively overcoming the two limitations identified in existing mixup methods. Empirical results from a variety of experimental scenarios indicate that SynerMix outperforms the best of either MixUp or SynerMix-Intra in accuracy by margins ranging from 0.1% to 3.43% (averaging 1.16%). It also surpasses the superior of either Manifold MixUp or SynerMix-Intra by margins ranging from 0.12% to 5.16% (averaging 1.11%).
SynerMix is versatile and can be applied to other research areas and tasks, such as speech classification, text classification, and object detection. It also paves the way for new potential research directions, which include: a) developing methods for predictive synthesis of feature representations, which leverage the shifting trajectories of image feature representations across epochs to accelerate model convergence, and b) investigating how the randomness of synthesized feature representations within local feature spaces affects the stochasticity of the gradient, with the aim of enhancing classification accuracy.